i need help on how to do these
1) x^2+1=3
2) x+4+5+1
3) 6x=12
please help

Answer by kekegeter (97)

1) x^2+1=3
x^2=2
x= 2 1/2
x= -2 1/2
The answer is: 2 1/2, -2 1/2
2) x+4+5+1
The answer is: x+10
3) 6x=12
6x/6=12/6
x=2^2*3/ 2*3
x=2^2-1
The answer is: x=2